# Latent Space Imaging

### v1

@@ 7 Data Availability
Figure S1: (a) Overview of all the blocks involved in the digital encoder architecture. The three steps capture coarse, middle, and fine details, with each step depending on the output of the previous one. (b) Illustration of the MIX process. The latent vector is statically projected to capture interactions among different levels of detail.
@@
The original StyleGAN2 implementation receives a random noise vector $\textit{Z}\in\mathbb{R}^{512}$ and utilizes a mapping network to project this to different levels of details, result in $\textbf{L}\in\mathbb{R}^{512\times l}$, $l$ may vary depending the desired image resolution that the generative models was trained to generate. In the case of face images, $l$ is equal to 18, finally we want O to map $\textit{J}\in\mathbb{R}^{d}$ to $\textbf{L}\in\mathbb{R}^{512\times 18}$. $d$ corresponds to our measurement count, therefore, going from 512 to 64.
@@
As depicted by Fig. S1a, the intermediate latent goes first through 3 different set of blocks from left to right. The first one produce the first 3 elements of L out of 18, each one of them only take into account the initial J. The middle block, besides J, it also takes the concatenation of $\textit{L}_{1},\textit{L}_{2}$ and $\textit{L}_{3}$ through the MIX module, as input, outputting $\textit{L}_{4}$ to $\textit{L}_{7}$. Finally, the last block follows the same idea, it concatenates the output of the middle one and pass through another MIX block, the output of this is summed together with the original measurement and it passes through the block to finally result in $\textit{L}_{8}$ to $\textit{L}_{18}$. The idea here is to make the optical-aware inversion network match the multi-level structure of StyleGAN2, following a coarse-to-fine approach.
@@
In the end, the latents from all detail levels pass through the MIX block. The purpose of the MIX blocks is to facilitate interactions and learn a weighted mixture across multiple detail levels. The model incorporates the spatial gating unit proposed in [49]. Within the MIX block (see Fig.S1b), the features are projected to a higher dimension, after which the resulting tensor ${x}$ is split into two parts, ${u}$ and ${v}$, along the channel dimension. This division allows for separate processing paths within the block. ${v}$ is normalized and linearly projected using a weights matrix and bias vector to capture interactions in a static manner, as the weights matrix remains unchanged after training because it does not depend on the input. The final element-wise multiplication of ${u}$ and the projected ${v}$ modulates the information flow, controlling which parts of ${u}$ are allowed to pass through—similar to a gating mechanism. Finally, this gated output is merged with the input through a residual connection, preserving the original information and enhancing gradient flow during back-propagation.
S2 Loss Functions and Training Details
@@
In this work, we trained the optical O and digital $\mathcal{D}_{\theta}$ encoders while keeping the generative model, StyleGAN2, frozen. To jointly optimize these components, we utilized an off-the-shelf inversion network pre-trained to invert images to the StyleGAN2 latent space of a specific domain [24]. This network is used to compute a latent similarity loss function ($\mathcal{L}_{lat}$), making the output of the $\mathcal{D}_{\theta}$ similar to the latent space approximated by this model. This process is described in Equation 3 of the main text.
@@
Additionally, we aim to match the image quality at the pixel level and enforce identity similarity. For this task, we used the identity loss ($\mathcal{L}_{id}$), which computes the cosine distance between feature maps extracted by ArcFace [50], a facial recognition network. We also utilized the $\ell_{2}$ norm ($\mathcal{L}_{l2}$) and LPIPS ($\mathcal{L}_{pips}$[51] to enforce pixel-wise and perceptual similarities, respectively.
@@
$\mathcal{L}_{total}=\lambda_{lat}\mathcal{L}_{lat}+\lambda_{id}\mathcal{L}_{id}+\lambda_{pips}\mathcal{L}_{pips}+\lambda_{l2}\mathcal{L}_{l2}$ (S1)
@@
$\begin{matrix}\mathcal{L}_{energy}=\frac{1}{d}\sum_{j}^{d}\left|\sum_{i}^{mn}\mathcal{\textbf{O}}_{i,j}-\epsilon_{j}\right|\end{matrix}$ (S2)
Here, $\epsilon$ represents the ground-truth energy level, heuristically designed to correspond to a certain percentage of pixels set to one. This percentage increases incrementally from $10\%$ to $90\%$, with a step of $1\%$. The patterns are shuffled to avoid any undesired structure.
To optimize O and $\mathcal{D}_{\theta}$, we employ distinct optimizers: Lion [52] for O and Ranger (combination of Lookahead [53] and the Rectified Adam [54]) for $\mathcal{D}_{\theta}$, each with a learning rate of $10^{-4}$. The weighting coefficients $\lambda$ for each loss function are set at $1,0.5,0.8,$ and $1$, in the respective order that they appear in Equation S1. Additionally, for the energy loss component, we set $\lambda$ to 3. Regarding batch sizes, we typically use 32 for face datasets. However, for the AFHQ [55] dataset, as discussed in Sec. S4, the batch size is reduced to 8 due to the limited number of images available. For all experiments we utilize a single NVIDIA A100.
S3 Experimental setup
@@
To enhance the congruence between our measurements and simulations, we employ a white image to determine a global scaling deviation and apply a correction factor. This approach effectively reduces the gap between simulation and measurement, yet some inherent challenges persist due to non-linear behaviors, minor deviations in DMD reflection angles, and sensor noise, which complicate the precise replication of simulated conditions. However, by leveraging the controlled environment of our experimental setup, we are able to systematically gather real-world training data and fine-tune $\mathcal{D}_{\theta}$.
@@
The fine-tuning process involves selecting a subset of 200 training images and measuring their outputs using our configured setup. We employ the loss functions specified in Equation S1, while reducing the learning rate to $10^{-5}$, to optimize performance and accuracy in real-world applications. Fig. S2 illustrates the effect of fine-tuning.
Figure S2: Demonstrates the effect of the fine-tuning process on our model. Post fine-tuning, the average ID similarity of the 100 test images increases from $0.25$ to $0.31$. This enhancement brings the values in line with those observed in our simulation results, indicating a closer match and improved performance.
S4 Additional Results
@@
We present additional LSI results encompassing a diverse set of images, along with the range of compression levels studied in this work. Our methodology involves reconstructing images with a resolution of $256\times 256$ pixels and examining compression rates that result in retaining 512, 256, 128, and 64 total measurements.
@@
In Fig. S3 a consistent retention of facial details is observed even with increasing compression, highlighting the effectiveness of our approach even at higher compression rates. Targeting the latent space addresses the over-smooth appearance typical of strong sub-sampling strategies, preserving textures and achieving a realistic look instead of creating flat reconstructions. Here, we emphasize the persistence of facial expressions, with key features such as eyebrow shape, beard, and smile being well reconstructed.
@@
Additional domains were explored in Fig. S4, we utilize our simulated pipeline with 512 measurements to reconstruct cats and dogs images from AFHQ dataset [55]. These datasets imposed additional challenge because they are very small compared with FFHQ, with only 5000 training images. However, LSI is capable to faithfully reconstruct such domain pictures.
Figure S4: Illustrates another set of reconstructions from our simulated results, highlighting the versatility of our methods across various domains, such as cats and dogs.
In Fig. S5 we also expand the examples of reconstructions performed by our physical setup utilizing the single-pixel approach for the 256 measurements case.
Figure S5: Showcase another set of reconstructions performed by the proposed physical setup.
S5 Failure Cases and Limitations
As we are focusing on faces and our underlying generative model is primarily trained with frontal-centered faces under typical daily lighting, strong deviations from these conditions are not properly modeled and result in strong hallucinations. Fig. S6a shows some failure conditions. From these observations, we note that accessories like glasses and hats are not modeled by the generative model and cannot be reconstructed. Additionally, other parts of the body are also not reproducible, and faces oriented far from a frontal view add additional burden to the reconstruction process. Similar behavior is observed in other works utilizing an encoder network to perform inversion when a purely software-based network is employed. In our case, we are dealing with a more challenging scenario where a real spatial encoder compresses the image to a very limited amount of data. Fig. S6b demonstrates how significant features can be lost in challenging scenes when compressing to over $99.9\%$. We notice that the mouth is not open as in the original image. This behavior can be explained by the underlying distribution learned by the generative model, where few images have the mouth in such a position. Consequently, our model struggles to reconstruct it accurately.
Refer to caption a b
Figure S6: (a) Illustrates the conditions where LSI fails to reproduce the scene. Demonstrate different challenging scenarios like accessories, face orientation, and color reproduction. (b) Exemplify a case where extreme compression hurt the expression reconstruction.
S6 Latent Space Image in the context of Compressive Sensing
Latent Space Imaging (LSI) aims to retrieve rich compressed information from a specific domain, directly targeting the latent space of the underlying reconstruction model. The concept can be used in different imaging modalities. Here, we showcase the compression aspect of LSI utilizing a single-pixel approach. Therefore, we compare against traditional compressive sensing (CS) strategies.
First, it is essential to state that some CS ideas in the literature are introduced only in simulation and may make unrealistic assumptions that prevent actual hardware implementation. Works like [56, 57, 58, 59], to name a few, provide valuable insights on how to improve compressive sensing reconstructions. However, they are not applicable because they have measurement matrices that are not feasible in the real world due to negative entries, or they do not consider challenges like quantization. Here, we are interested in methods that are feasible in an experimental setup.
In this context, DCAN [60] and follow-up works optimize their binary patterns together with a deep learning model for reconstruction in the spatial domain. Other solutions shift to the Fourier domain as proposed by [18], leveraging the fact that information is usually concentrated on low-frequencies, enabling under-sampling strategies with high-quality reconstructions but suffering when dealing with low sampling rates [61]. Recent Fourier single-pixel imaging (FSI) techniques try to overcome such issue, using a more advanced network as a reconstruction method, improving the heuristics for the frequency domain filters designing and jointly optimizing the masks and reconstruction algorithms [25, 62, 63].
Finally, to assess how LSI reconstruction quality is compared with SPI, we retrain FSI models using a fixed circle mask and optimize it with the deep learning reconstruction using AuSamNet pipeline [25]. To also analyze the performance when utilizing more sophisticated deep learning strategies from a CS perspective, we retrain SAUNet [59] imposing the quantization constraints. We train and evaluate using the same dataset as LSI and a similar compression rate.
@@
When evaluating under our heavily compressed setting, AuSamNet cannot optimize the mask, leading to unstable results. FSI-DL is capable of reconstructing the images but produces low-quality results as demonstrated by Fig. S7.
@@
Similar to FSI-DL, SAUNet can reconstruct images but struggles to maintain fine details, resulting in lower quality outputs (see Fig. S7). It is important to note that we trained SAUNet following the methodology outlined in their work. Like many other compressive sensing approaches, we utilized the YCrCb color space. In this approach, the objective is to reproduce the luminance component, while the chrominance components were added during inference using ground-truth values. This approach gives them an advantage, as they would in any practical setting require additional measurements to reproduce color images or model color filters in conjunction with the modulation masks, as proposed by AuSamNet [25].
Figure S7: LSI qualitatively compared against compressive sensing approaches across several compression ratios.
S6.1 Different Generative Models
Recent advancements in diffusion models have significantly improved the generation of high-quality, text-guided images, their success root also in leveraging the latent space [64]. This process utilizes a VAE as the encoder to the latent space and decoder to back to image, which can also be applied in LSI. However, such generative models still depend on spatial structures, with a latent space of $4\times 32\times 32$ for $256\times 256$ images. To leverage the encoding power of VAE, we replaced the StyleGAN generator with the VAE decoder and used the encoder to provide a new ground-truth $\mathcal{E}(\textit{I})$ for our target latent space, then solving the Equation 3 in the main text. The spatial structure in this latent space, coupled with its training for multi-domain applications, leads to some traditional compressive sensing issues, where high-frequency information is often lost during reconstruction (see Fig.3c in the main text). However, because our approach directly images to a compressed latent space, it outperforms other methods like Fourier single-pixel imaging. The compactness of the StyleGAN latent space exemplifies its superior suitability over VAE for the application demonstrated here.
S7 Evaluation Metrics
It is well established that traditional metrics such as PSNR and SSIM do not correlate well with human visual perception [51, 65, 66]. These metrics prioritize pixel-level accuracy and are largely insensitive to aspects of perceptual quality, failing to account for high-level structural information. Seminal studies, such as [51], have demonstrated that phenomena like blurring, which significantly impact perceptual experience, often lead to minimal changes in the L2 norm. This limitation becomes particularly pronounced in the context of heavily compressed image reconstructions, where PSNR and SSIM inadequately reflect the extent of perceived quality deterioration.
In this work, we evaluate our methods in terms of identity similarity and FID [67]. The first one, is the dot product between features extracted from CurricularFace [26], a specialized model for face recognition, which is different from the one utilized during training to compute the ID loss, in order to guarantee that we are not optimizing for specific evaluation metric.
Finally, Table S1 presents the evaluation metrics for LSI, FSI-DL, and SAUNet. It is evident that our method produces superior perceptual quality, particularly in heavily compressed cases where other methods fail entirely.
Table S1: Comparison of ID and FID scores across different methods and resolutions (top result shown in bold).
512 256 128 64
ID $\uparrow$ FID $\downarrow$ ID $\uparrow$ FID $\downarrow$ ID $\uparrow$ FID $\downarrow$ ID $\uparrow$ FID $\downarrow$
LSI 0.32 35.57 0.31 35.65 0.27 37.19 0.22 37.29
FSI-DL 0.12 107.48 0.07 118.79 0.04 134.40 0.02 173.71
SAUNet 0.24 58.96 0.14 75.53 0.09 104.23 0.06 107.35

### v2

@@ B.1 Digital Encoder Architecture
Figure B.1: (Top) Overview of all the blocks involved in the digital encoder architecture. The three steps capture coarse, middle, and fine details, with each step depending on the output of the previous one. (Bottom) Illustration of the MIX process. The latent vector is statically projected to capture interactions among different levels of detail.
@@
The original StyleGANXL implementation receives a random noise vector and utilizes a mapping network to project this to different levels of details, result in $\textbf{L}\in\mathbb{R}^{512\times l}$, $l$ may vary depending the desired image resolution that the generative models was trained to generate. In the case of face images, $l$ is equal to 18, finally we want O to map $\textit{J}\in\mathbb{R}^{d}$ to $\textbf{L}\in\mathbb{R}^{512\times 18}$. $d$ corresponds to our measurement count, therefore, going from 512 to 4.
@@
As depicted by Fig. B.1, the intermediate latent goes first through 3 different set of blocks from left to right. The first one produces the first 3 elements of L out of 18, each one of them only take into account the initial J. The middle block, besides J, it also takes the concatenation of $\textit{L}_{1},\textit{L}_{2}$ and $\textit{L}_{3}$ through the MIX module, as input, outputting $\textit{L}_{4}$ to $\textit{L}_{7}$. Finally, the last block follows the same idea, it concatenates the output of the middle one and pass through another MIX block, the output of this is summed together with the original measurement and it passes through the block to finally result in $\textit{L}_{8}$ to $\textit{L}_{18}$. The idea here is to make the optical-aware inversion network match the multi-level structure of StyleGANXL, following a coarse-to-fine approach.
@@
In the end, the latents from all detail levels pass through the MIX block. The purpose of the MIX blocks is to facilitate interactions and learn a weighted mixture across multiple detail levels. The model incorporates the spatial gating unit proposed in [41]. Within the MIX block (see Fig.B.1), the features are projected to a higher dimension, after which the resulting tensor ${x}$ is split into two parts, ${u}$ and ${v}$, along the channel dimension. This division allows for separate processing paths within the block. ${v}$ is normalized and linearly projected using a weights matrix and bias vector to capture interactions in a static manner, as the weights matrix remains unchanged after training because it does not depend on the input. The final element-wise multiplication of ${u}$ and the projected ${v}$ modulates the information flow, controlling which parts of ${u}$ are allowed to pass through—similar to a gating mechanism. Finally, this gated output is merged with the input through a residual connection, preserving the original information and enhancing gradient flow during back-propagation.
B.2 Loss Functions and Training Details
@@
In this work, we trained the optical O and digital $\mathcal{D}_{\theta}$ encoders while keeping the generative model, StyleGANXL, frozen. To jointly optimize these components, we trained an off-the-shelf inversion network pre-trained to invert images to the latent space of a specific domain [51]. This network is used to compute a latent similarity loss function ($\mathcal{L}_{lat}$), making the output of the $\mathcal{D}_{\theta}$ similar to the latent space approximated by this model. This process is described in Equation 3 of the main text.
@@
Additionally, we aim to match the image quality at the pixel level and enforce identity similarity. For this task, we used the identity loss ($\mathcal{L}_{id}$), which computes the cosine distance between feature maps extracted by ArcFace [13], a facial recognition network. We also utilized the $\ell_{2}$ norm ($\mathcal{L}_{l2}$) and DINO/LPIPs features loss($\mathcal{L}_{p}$[79, 76] to enforce pixel-wise and perceptual similarities, respectively. The details of DINO feature extraction are provided in [79].
@@
$\mathcal{L}_{total}=\lambda_{lat}\mathcal{L}_{lat}+\lambda_{id}\mathcal{L}_{id}+\lambda_{p}\mathcal{L}_{p}+\lambda_{l2}\mathcal{L}_{l2}$ (5)
@@
$\begin{matrix}\mathcal{L}_{energy}=\frac{1}{d}\sum_{j}^{d}\left|\sum_{i}^{mn}\mathcal{\textbf{O}}_{i,j}-\epsilon_{j}\right|\end{matrix}$ (6)
Here, $\epsilon$ represents the ground-truth energy level, heuristically designed to correspond to a certain percentage of pixels set to one. This percentage increases incrementally from $10\%$ to $90\%$, with a step of $1\%$. For strong compression from 1:1024 to 1:16384 the min and max energy boundaries move closer to $50\%$ accordingly. The patterns are shuffled to avoid any undesired structure. As demonstrated in Fig. B.2, with the energy loss applied, the pixel occupancy histogram showcases a broad spectrum of occupancies, reflecting a high diversity of patterns. In contrast, without this modification, the patterns are more uniform in their occupancy levels, exhibiting less variability.
Figure B.2: Shows the histogram of pixel occupancy, comparing scenarios where the energy loss function is and is not utilized. The use of energy modeling enhances the dynamic range of the measurements captured by the single sensor.
To optimize O and $\mathcal{D}_{\theta}$, we employ distinct optimizers: Lion [7] for O and Ranger (combination of Lookahead [75] and the Rectified Adam [42]) for $\mathcal{D}_{\theta}$, each with a learning rate of $10^{-4}$. The weighting coefficients $\lambda$ for each loss function are set at $1,0.5,0.8,$ and $1$, in the respective order that they appear in Equation 5. Additionally, for the energy loss component, we set $\lambda$ to 3. Regarding batch sizes, we typically use 32 for face datasets. However, for the AFHQ [8] dataset, as discussed in Sec. C, the batch size is reduced to 8 due to the limited number of images available. For all experiments we utilize a single NVIDIA A100.
B.3 Optical Encoder Optimization Details
In the O optimization process, in order to deal with the quantized patterns, we employ the straight-through estimator (STE) [23], a technique also favored by VQ-VAEs [62, 50, 43]. During the forward pass, values are quantized and constrained before computing J. However, during back-propagation, gradients are allowed to flow through the non-quantized version, facilitating weight updates. Importantly, we ensure that the entries of O remain positive and are bounded between $[0,1]$. Although using complementary patterns [72] allows for the inclusion of negative values, empirical experiments have shown that this approach does not yield significant improvements and necessitates double the measurements.
Initially, masks are generated from a uniform distribution and are then binarized through a quantization process. This process results in binary patterns that achieve an even total counting of 0s and 1s, theoretically maximizing bit entropy and steering the system towards an optimal solution [39]. This distribution presumes the sensor is sensitive enough to discern very fine differences, a demanding prerequisite as each pattern will have similar total intensity. This issue was addressed using the energy loss mechanism.
B.4 Experimental setup
Figure B.3: Experiment setup. An objective lens collects light from monitor and relays it to the DMD, which modulates the light with reflecting patterns. A relay lens directs the modulated light to the photodiode, and the PC records the measurements. Yellow arrows denote the light path.
@@ B.4 Experimental setup
To enhance the congruence between our measurements and simulations, we employed a white image to determine a global scaling deviation and applied a correction factor. This approach effectively reduces the gap between simulation and measurement, yet some inherent challenges persist due to non-linear behaviors, minor deviations in DMD reflection angles, and sensor noise, which complicate the precise replication of simulated conditions. However, leveraging the controlled environment of our experimental setup, we are able to systematically gather real-world training data and fine-tune $\mathcal{D}_{\theta}$.
@@
The fine-tuning process involves selecting a subset of 200 training images and measuring their outputs using our configured setup. We employ the loss functions specified in Equation 5, while reducing the learning rate to $10^{-5}$, to optimize performance and accuracy in real-world applications.
C Part III: Additional Results
@@
We present additional LSI results encompassing a diverse set of images, along with the range of compression levels.
@@
Figures C.7C.8, and C.9 demonstrate a consistent retention of facial details, particularly as compression increases, highlighting the effectiveness of our approach at higher compression rates. By targeting the latent space, our method addresses the over-smoothing often observed with aggressive sub-sampling strategies, preserving textures and producing realistic reconstructions instead of flat, featureless outputs. Notably, key facial features such as eyebrow shape, beards, and smiles are well reconstructed, emphasizing the persistence of facial expressions. In contrast, competitor models fail to achieve similar results.
For the downstream tasks, Fig. C.6 presents additional results from the experimental setup for attribute classification. To further evaluate and enhance the assessment of our method, we include an extensive set of qualitative comparisons (see Figs. C.10C.11C.12C.13C.14, and C.15) against other methods. These comparisons focus on simulated results for face segmentation and landmark detection.
Table C.1 quantitatively summarizes our results, highlighting the overall superiority of our method in reconstruction and downstream tasks, particularly in high-compression setups.
Additional domains were explored in Fig. C.4, we utilize our simulated pipeline with 512 (1:128) measurements to reconstruct cats and dogs images from AFHQ dataset [8]. These datasets imposed additional challenge because they are very small compared with FFHQ, with only 5000 training images. However, LSI is capable to faithfully reconstruct such domain pictures with a correspondingly trained encoder.
Figure C.4: Illustrates another set of reconstructions from our simulated results, highlighting the versatility of our methods across various domains, such as cats and dogs. For this dataset we utilized StyleGAN2 instead XL, also showing the versatility of our method utilizing different generative models.
Figure C.5: Illustrates the conditions where LSI fails to reproduce the scene since the target scene is out of distribution of the generative model.
Figure C.6: Additional attribute visualizations from the experimental validation are presented. The leftmost column displays the original faces as shown on the monitor, followed by reconstructed images at compression ratios of 1:512 and 1:2048, which correspond to 128 and 32 measurements, respectively.
C.1 Out-of-Distribution Cases and Limitations
As our focus is on faces, and our underlying generative model is primarily trained on frontal-centered faces under typical daily lighting conditions, significant deviations from these conditions are not well-modeled and often result in noticeable hallucinations. Fig. C.5 illustrates several out-of-domain scenarios. For instance, providing a non-facial input, such as a cat, to the face generative model results in a hallucinated face that aligns with the characteristics of the given input signal. Similarly, accessories like hats or heavy makeup cannot be accurately modeled, as such data lies outside the model’s training distribution.
C.2 Qualitative and Quantitative Comparisons
We evaluated a diverse set of representative models encompassing various approaches, including block-based sensing matrices, unfolding networks, single-pixel imaging techniques, and deep learning-driven single-pixel frameworks. Additionally, we incorporated optimization-based methods that leverage generative models as priors. To provide a broader perspective, we introduced a variant of our pipeline, referred to as non-latent, which replaces the generative model and latent representation with a state-of-the-art reconstruction network and a conventional signal recovery method. Further details on this model are provided below.
Preliminary notes on comparisons:
  • All methods were retrained on the same dataset and evaluated on the same testing dataset as LSI.
  • SAUNet, FSI-DL, and Non-latent LSI utilize a quantized (binary) sensing matrix, similar to LSI. In contrast, the sensing matrices of other methods can represent any float value, giving them an (unrealistic!) advantage.
  • SAUNet, OCTUF [58], and TCS-Net [17] are trained to reproduce luminance only, as proposed in their original papers. During inference, chrominance components were added using ground-truth values. Other methods, however, consider color images, modeling them either implicitly (via latent space) or explicitly (via demosaicking).
  • For comparison purposes, downstream applications for all competitors were evaluated using FaRL [78], based on their reconstructed images, as they lack the inherent capability to perform these tasks. Because of low reconstruction quality, some competitor methods may fail the downstream tasks, we discard such data points only for the specific method. Notably, for LSI, these results were achieved through a simple linear projection, entirely eliminating the need for a separate, complex model tailored to each specific task.
@@ C.2 Qualitative and Quantitative Comparisons
LSI 1:128 91.97% 92.74% 27.38 89.07% 70.00% 1.48
1:256 90.98% 92.68% 26.62 89.15% 70.94% 1.43
1:512 89.61% 91.67% 28.66 89.20% 70.25% 1.48
1:1024 81.12% 87.44% 28.79 88.74% 69.18% 1.52
SAUNet 1:128 39.47% 64.12% 58.96 83.05% 74.15% 1.36
1:256 17.84% 50.12% 75.53 82.66% 71.32% 1.59
1:512 5.52% 31.42% 104.23 81.08% 66.79% 1.99
1:1024 2.86% 16.59% 107.35 79.98% 62.37% 2.20
FSI-DL 1:128 15.28% 32.60% 107.48 81.66% 62.20% 2.00
1:256 3.30% 13.16% 118.79 79.43% 53.38% 3.20
1:512 1.82% 1.93% 134.40 76.72% 41.45% 5.06
1:1024 0.85% 2.05% 173.71 75.96% 34.95% 6.06
OCTUF 1:128 54.99% 63.54% 70.83 83.93% 75.78% 1.31
1:256 59.27% 73.09% 77.16 83.99% 74.78% 1.26
1:512 16.63% 43.31% 96.17 82.33% 69.92% 1.73
1:1024 5.83% 26.62% 99.85 81.13% 66.18% 1.99
TCS-Net 1:128 25.15% 49.69% 119.42 82.21% 66.00% 1.71
1:256 1.32% 5.50% 248.60 77.71% 52.11% 3.31
1:512 1.07% 2.62% 296.48 75.19% N/A 7.43
1:1024 0.11% 1.62% 333.34 74.05% N/A 15.07
CSGM 1:128 20.04% 51.85% 44.46 82.17% 62.20% 1.89
1:256 13.66% 39.16% 46.30 81.28% 56.96% 2.36
1:512 5.87% 21.96% 53.12 78.68% 46.07% 3.29
1:1024 2.00% 10.23% 63.70 76.16% 38.31% 4.15
Non-Latent 1:128 94.22% 94.84% 72.96 84.47% 77.47% 1.07
1:256 87.52% 91.49% 75.62 83.71% 74.59% 1.26
1:512 71.82% 82.98% 81.87 83.08% 70.10% 1.53
1:1024 41.47% 68.22% 89.04 82.09% 64.73% 1.90
Table C.1: Latent Space Imaging (LSI) and competing methods were quantitatively evaluated on simulated results across various downstream tasks, with compression ratios ranging from 1:128 to 1:1024. The results emphasize the superior performance of LSI, particularly in highly compressed scenarios. For the image reconstruction task, evaluations were conducted using the VGGFace [12] and DLib [32] image recognition pipelines. Metrics assessed include Fréchet Inception Distance (FID), classification accuracy, F1-mean score for segmentation, and Normalized Mean Error (NME), normalized by the diagonal of the face bounding box. Bold numbers highlight the best model for each metric and setting.
@@
When evaluating under our heavily compressed setting, AuSamNet cannot optimize the mask, leading to unstable results. FSI-DL is capable of reconstructing the images but produces low-quality results as demonstrated by Figures C.7, C.8, C.9.
@@
OCTUF [58] and TCS-Net [17]
Both methods utilize transformer-based reconstruction techniques integrated with block-based image compressed sensing for measurements, where the image is divided into smaller patches. However, the sensing matrices used are not restricted to values below 1, making their physical implementation unfeasible. The methods were trained using the original code without any modifications, apart from adjustments to the compression ratios. Finally, they benefit from an unbounded sensing matrix and similar to SAUNet [64], they do not model color, only learning the luminance.
Optimization-Based (CSGM [24])
leverages a generative model as a prior to reconstruct compressed signals. However, its measurement matrix is neither bounded nor quantized, making it unsuitable for physical deployment. Additionally, it requires multiple iterations and cannot achieve reconstruction with a single forward pass.
Non-Latent LSI
We further evaluate our approach by implementing a version of LSI that does not utilize the latent space. In this configuration, the 2D image $\in\mathbb{R}^{256\times 256}$ is reconstructed directly from the measurements using Differential Ghost Imaging (DGI) [16], followed by a state-of-the-art image reconstruction model [9], similar to the methodology described by [63]. This replaces the latent space representation and generative model entirely with a direct reconstruction pipeline. The training process, including perceptual and identity loss functions, Optical Encoder (O) optimization, and other parameters, remain unchanged. While this version performs well for lower compression ratios (1:128), it is consistently outperformed by LSI, particularly at higher compression levels such as 1:512 and 1:1024 (see Table C.1), underscoring the critical role of latent space representation.
[…]
LSI Random 48.9% 71.7% 37.20
LSI 90.98% 92.68% 26.62
Table C.2: Quantitative comparison between LSI and LSI with fixed random patterns (akin to classical Compressed Sensing), emphasizing the critical role of jointly optimizing the optical encoder for enhancing the performance. Compression ratio of 1:256.
C.3 Background Discussion
To emphasize facial features, we mask out the background to prevent our already highly compressed signal from being used on non-facial information. It is worth noting, however, that our method remains effective even when trained on faces with uncontrolled backgrounds, achieving identity classification accuracies of $86.26\%$ and $86.88\%$ with VGGFace and DLib, respectively. Ablation study conducted for a compression ratio of 1:256.
C.4 Optical Encoding vs. Fixed Random Encoding
Optimizing the optical encoding plays a critical role in image reconstruction. This is evident in Table C.2, which shows a significant performance drop when LSI is trained using fixed random patterns instead optimized encoding towards the latent space.
Figure C.7: Different methods compared for facial reconstruction quality.
Figure C.8: Different methods compared for facial reconstruction quality.
Figure C.9: Different methods compared for facial reconstruction quality.
Figure C.10: Facial segmentation comparison among different methods.
Figure C.11: Facial segmentation comparison among different methods.
Figure C.12: Facial segmentation comparison among different methods.
Figure C.13: Landmarks detection comparison among different methods.
Figure C.14: Landmarks detection comparison among different methods.
Figure C.15: Landmarks detection comparison among different methods.